# Explicit Performance Bound of Finite Blocklength Coded MIMO: Time-Domain versus Spatiotemporal Channel Coding

### v2

@@ Abstract
In the sixth generation (6G), ultra-reliable low-latency communications (URLLC) will further develop to achieve $\text{TK}\upmu$ extreme connectivity, and multiple-input multiple-output (MIMO) is expected to be a key enabler for its realization. On the premise of ensuring the same rate and reliability, the spatial domain advantage of MIMO has the potential to further shorten the time-domain code length. Different coded MIMO schemes exhibit disparities in exploiting the spatial domain characteristics, so it is necessary to theoretically elucidate the performance of different coding schemes to illustrate how MIMO can support 6G $\text{TK}\upmu$ transmission. We consider two extreme MIMO coding schemes, namely, time-domain coded MIMO where the codewords on multiple spatial channels are independent of each other, and spatiotemporal coded MIMO in which multiple spatial channels are jointly coded. In this paper, by analyzing the statistical characteristics of information density and utilizing the normal approximation, we provide explicit performance bounds for finite blocklength coded MIMO under time-domain coding and spatiotemporal coding. We find that spatiotemporal coding can effectively alleviate the performance degradation induced by short blocklength by improving the spatial degrees of freedom (DoF). However, for time-domain coding, the independent coding of each spatial link results in more severe performance degradation due to shorter blocklength, even when a relatively large spatial DoF is available. These results indicate that spatiotemporal coding can optimally exploit the spatial dimension advantages of MIMO systems compared with time-domain coding, enabling extremely low error-rate communication under stringent blocklength constraint.
@@
MIMO, URLLC, finite blocklength, time-domain coding, spatiotemporal coding, explicit performance bound.
@@ I Introduction
Since URLLC relies on shorter packet length and smaller transmission interval [6], the length of codewords will be substantially reduced to meet the stringent delay constraint. To meet the continuously evolving demands of URLLC, finite blocklength coding addresses the practical latency constraints by optimizing the trade-off between blocklength (latency), decoding error probability (reliability), and coding rate [7]. In [8], the nonasymptotic upper and lower bounds and normal approximations on coding rate as a function of decoding error probability and blocklength at finite blocklength for the binary erasure channel (BEC), the binary symmetric channel (BSC), and the additive white Gaussian noise (AWGN) channel are summarized and analyzed. Concurrently, the specific derivation of the normal approximation in the AWGN channel is presented in [9]. In a succession of subsequent research efforts, these results have been extended to other types of scalar channels, such as [10, 11, 12, 13]. Obviously, in scalar channels, the amount of data information decreases with decreasing blocklength for given decoding error probability, and it evidently fails to meet the progressively increasing demands of 6G URLLC.
@@
Multiple-input multiple-output (MIMO) technology introduces additional spatial domain degree of freedom (DoF) by deploying multiple antennas at both ends of the transceivers [14]. MIMO is considered to be a key technology for guaranteeing communication performance under low latency requirements in URLLC scenarios [15, 16, 17]. There have been a series of research efforts to extend the normal approximation in [8] to MIMO systems. Despite the fact that the normal approximation is not as accurate as the saddlepoint method [18, 19, 20] at low coding rates, its concise form makes it easier to intuitively analyze the relationship between various key performance indicators and theoretically illustrate the internal logic of MIMO supporting URLLC transmission. In [21] and [22], the coding rate of a quasi-static MIMO channel was described given blocklength and decoding error probability. [23] presented the closed-form expression of the channel dispersion in MIMO coherent fading channel, and [24] presented a high-SNR normal approximation of the maximum coding rate in a non-coherent Rayleigh block-fading MIMO channel. Meanwhile, [25] employed the information spectrum method and random matrix theory to derive the second-order coding rate of the MIMO quasi-static Rayleigh fading channel. Given that the rate-latency-reliability relationship obtained from the aforementioned research is overly complicated to directly infer the effect of the number of antennas on the performance of a finite blocklength MIMO system, [26] further derived the explicit solutions for the average maximal achievable rate in MIMO systems with respect to spatial DoF, decoding error probability and blocklength based on the conclusion of [21]. The results show that the spatial DoF (coding in spatial-temporal domain) can effectively compensate for the performance loss caused by the reduction of temporal DoF due to shortened codewords. [27] provided the theoretical upper bound for spatiotemporal two-dimensional (2-D) channel coding and suggested that through spatiotemporal coding, transmission reliability and latency can be flexibly balanced in a variety of communication scenarios including URLLC. In fact, we can find that the above finite blocklength theoretical studies on MIMO channels all adopt the joint spatiotemporal coding scheme with multiple antennas at the transmitter. However, the time-domain coding scheme with independent coding at each transmit antenna can also be used in practical MIMO coding. A spatiotemporal 2-D polar coding scheme was proposed in [28], and simulation analysis reveals that spatiotemporal coding is capable of fully exploiting the spatial domain resources to achieve high data-rate performance under low latency constraint, while time-domain coding will not benefit from spatial resources in comparison.
@@
To the best of our knowledge, the intuitive performance superiority of spatiotemporal coding over time-domain coding has not yet been thoroughly investigated from a theoretical point of view. In order to comprehensively understand the benefits of spatiotemporal coding for URLLC transmission in MIMO systems, it is essential to derive explicit solutions for the achievable rates of MIMO systems employing spatiotemporal and time-domain coding schemes. The theoretical tools in [8, 9, 21, 22, 23, 24, 25, 26, 27] enable one to characterize the maximal achievable rates of spatiotemporal coding and time-domain coding by analytical means, and it is of great significance to additionally find an explicit expression to identify how spatial DoF improves overall performance. Furthermore, novel theoretical studies in finite blocklength coding have the potential to provide a new perspective into practical spatiotemporal coding design.
@@
The remainder of the paper is organized as follows. In Section II, we give the specification of information density and binary hypothesis testing, based on which the definitions of achievability and converse bounds are provided. The explicit performance bounds of finite blocklength coded MIMO in different coding modes are derived in Section III. Section IV analyzes the normalized maximal achievable rate and decoding error probability in different coding modes. Section V provides numerical results to reveal the advantages of spatiotemporal coding versus time-domain coding in finite blocklength coded MIMO.
@@ II Preliminaries
In this section, we introduce the definitions of information density and binary hypothesis testing under the memoryless channel model fixing the output distribution (such as AWGN channel), thereby presenting the expression of the Neyman-Pearson function. Based on this, we provide the achievability bound and converse bound to facilitate the subsequent solution of the normal approximations for different coding schemes of the MIMO channel.
@@ II-A Information Density
For a joint distribution ${{P}_{{\mathbf{X}}{\mathbf{Y}}}}$ on ${{\mathcal{A}}^{n}}\times{{\mathcal{B}}^{n}}$ and under the assumption of a memoryless channel, the information density is
$\iota\left({\mathbf{x}};{\mathbf{y}}\right)$ $=\log\frac{d{{P}_{{\mathbf{Y}}|{\mathbf{X}=\mathbf{x}}}}}{d{{P}_{{\mathbf{Y}}}}}\left({\mathbf{y}}\right)$
$=n\log\frac{d{{P}_{{{Y}}|{{X}={x}}}}}{d{{P}_{{{Y}}}}}\left({{y}}\right)=n\iota\left({{x}};{{y}}\right).$ (6)
@@ II-B Binary Hypothesis Testing
Proposition 1
Using the Berry-Esseen CLT, the Neyman-Pearson function can be expressed as
$\log{{\beta}_{\alpha}^{n}}\left({{P}_{{\mathbf{Y}}|{\mathbf{X}=\mathbf{x}}}},{{P}_{{\mathbf{Y}}}}\right)=-nC+\sqrt{nV}\Phi^{-1}\left(1-\alpha\right)+\mathcal{O}\left(\log n\right).$ (13)
where $C$ and $V$ are represented by the first and second moments of the information density, respectively,
$C=$ $\frac{1}{n}\mathbb{E}\left[\iota\left({\mathbf{x}};{\mathbf{Y}}\right)\right],$ (14)
$V=$ $\frac{1}{n}\text{Var}\left[\iota\left({\mathbf{x}};{\mathbf{Y}}\right)\right].$ (15)
Proof: See Appendix A.
@@ III Performance Bounds of MIMO Systems under Different Coding Modes
In this section, we first present the MIMO channel model and analyze the difference between spatiotemporal coding and time-domain coding. Subsequently, we proceed to solve the upper bounds of performance that MIMO channels can achieve in spatiotemporal coding and time-domain coding, respectively.
@@ III-A MIMO Channel Model
We consider a quasi-static flat Rayleigh fading MIMO channel, wherein the random fading coefficients remain invariant throughout the duration of each codeword. This is a typical assumption for URLLC where the blocklength is usually short enough. The relationship between channel input and channel output can be expressed as
@@
For channel coding in MIMO systems, two coding methods with different utilization of spatial DoF, spatiotemporal coding and time-domain coding, can be considered. The difference between the two methods is illustrated in Fig. 1. In the case of spatiotemporal coding, each bit stream is jointly encoded within both the time domain and the spatial domain, thereby generating a complete codeword $\mathsf{X}$. Then, the corresponding signals of the codeword in different spatial domains are sent out through $L$ transmit antennas. In the case of time domain coding, $L$ bit streams are encoded in the time domain. The codewords ${{\mathbf{x}}_{1}},\ldots,{{\mathbf{x}}_{L}}$ corresponding to each column of the resulting transmitted signal $\mathsf{X}$ are independent of each other. Thereafter, the time-domain codewords are dispatched through $L$ transmit antennas.
@@ III-B Spatiotemporal Coding
$=n\log\det\left({\mathsf{I}_{N}}+\frac{1}{n}{\mathsf{H}^{\text{H}}}\mathsf{Q}_{X}\mathsf{H}\right)$
$\quad+\text{tr}\left[{{{\left(\mathsf{X}\mathsf{H}+\mathbb{W}\right)}^{\text{H}}}\left(\mathsf{X}\mathsf{H}+\mathbb{W}\right)}\left({{\mathsf{I}_{N}}+\frac{1}{n}{\mathsf{H}^{\text{H}}}\mathsf{Q}_{X}\mathsf{H}}\right)^{-1}-{{\mathbb{W}}^{\text{H}}}\mathbb{W}\right]\log e$
$=n\log\det\left({\mathsf{I}_{N}}+\frac{1}{n}{\mathsf{H}^{\text{H}}}\mathsf{Q}_{X}\mathsf{H}\right)$
$\quad+\text{tr}\left[\left({-\frac{1}{n}{\mathsf{H}^{\text{H}}}\mathsf{Q}_{X}\mathsf{H}{{\mathbb{W}}^{\text{H}}}\mathbb{W}+{{\mathbb{W}}^{\text{H}}}\mathsf{X}\mathsf{H}+{\mathsf{H}^{\text{H}}}{\mathsf{X}^{\text{H}}}\mathbb{W}+\mathsf{H}^{\text{H}}{\mathsf{X}^{\text{H}}}\mathsf{X}\mathsf{H}}\right)\left({{\mathsf{I}_{N}}+\frac{1}{n}{\mathsf{H}^{\text{H}}}\mathsf{Q}_{X}\mathsf{H}}\right)^{-1}\right]\log{e},$ (21)
@@
Due to the spherical symmetry of both $\mathcal{F}^{\text{ST}}$ (equivalent to ${{P}_{{\mathbb{Y}}|{\mathbb{X}=\mathsf{X},\mathbb{H}=\mathsf{H}}}}$) and
@@
$\iota\left(\mathsf{X};\mathbb{Y}|\mathsf{H}\right)$ $=n\log\det\left({\mathsf{I}_{N}}+\frac{\rho}{L}{\mathsf{H}^{\text{H}}}\mathsf{H}\right)$
$\quad+\text{tr}\left[\left({-\frac{\rho}{L}{\mathsf{H}^{\text{H}}}\mathsf{H}{{\mathbb{W}}^{\text{H}}}\mathbb{W}+{{\mathbb{W}}^{\text{H}}}\mathsf{X}\mathsf{H}+{\mathsf{H}^{\text{H}}}{\mathsf{X}^{\text{H}}}\mathbb{W}+n\frac{\rho}{L}\mathsf{H}^{\text{H}}\mathsf{H}}\right)\left({{\mathsf{I}_{N}}+\frac{\rho}{L}{\mathsf{H}^{\text{H}}}\mathsf{H}}\right)^{-1}\right]\log{e}$
@@
$\quad+\text{tr}\left[n{\mathsf{I}_{N}}-\left({\frac{\rho}{L}{\mathsf{H}^{\text{H}}}\mathsf{H}{{\mathbb{W}}^{\text{H}}}\mathbb{W}-{{\mathbb{W}}^{\text{H}}}\mathsf{XH}-{\mathsf{H}^{\text{H}}}{\mathsf{X}^{\text{H}}}\mathbb{W}+n{\mathsf{I}_{N}}}\right)\left({{\mathsf{I}_{N}}+\frac{\rho}{L}{\mathsf{H}^{\text{H}}}\mathsf{H}}\right)^{-1}\right]\log e$
@@
where ${{W}_{i,j}}\sim\mathcal{CN}\left(0,1\right)$. In this case, according to Proposition 1, it can be derived that
${{\beta}_{1-\varepsilon}^{n}}\left({{P}_{{\mathbb{Y}}|{\mathbb{X}=\mathsf{X},\mathbb{H}=\mathsf{H}}}},{{P}_{{\mathbb{Y}}|\mathbb{H}=\mathsf{H}}}\right)=-n{{C}^{\text{ST}}}\left(\mathsf{H}\right)+\sqrt{n{{V}^{\text{ST}}}\left(\mathsf{H}\right)}\Phi^{-1}\left(\varepsilon\right)+\mathcal{O}\left(\log n\right),$ (23)
Subsequently, the channel capacity and channel dispersion of the MIMO channel with spatiotemporal codewords in a channel use are expressed as
@@
${{V}^{\text{ST}}}\left(\mathsf{H}\right)=\frac{1}{n}\text{Var}\left[\iota\left(\mathsf{X};\mathbb{Y}|\mathsf{H}\right)\right]=\sum\limits_{i=1}^{m}{\left[1-\frac{1}{{{\left(1+\frac{\rho}{L}{{\lambda}_{i}}\right)}^{2}}}\right]}{{\log}^{2}}e=\sum\limits_{i=1}^{m}{{{V}_{i}}},$ (25)
where

### v3

@@ Abstract
In the sixth generation (6G), ultra-reliable low-latency communications (URLLC) will be further developed to achieve $\text{TK}\upmu$ extreme connectivity. On the premise of ensuring the same rate and reliability, the spatial domain advantage of multiple-input multiple-output (MIMO) has the potential to further shorten the time-domain code length and is expected to be a key enabler for the realization of $\text{TK}\upmu$. Different coded MIMO schemes exhibit disparities in exploiting the spatial domain characteristics, so we consider two extreme MIMO coding schemes, namely, time-domain coding in which the codewords on multiple spatial channels are independent of each other, and spatiotemporal coding in which multiple spatial channels are jointly coded. By analyzing the statistical characteristics of information density and utilizing the normal approximation, we provide explicit performance bounds for finite blocklength coded MIMO under time-domain coding and spatiotemporal coding. It is found that, different from the phenomenon in time-domain coding where the performance declines as the blocklengths decrease, spatiotemporal coding can effectively compensate for the performance loss caused by short blocklengths by improving the spatial degrees of freedom (DoF). These results indicate that spatiotemporal coding can optimally exploit the spatial dimension advantages of MIMO systems, enabling extremely low error-rate communication under stringent blocklengths constraint.
@@
MIMO, URLLC, finite blocklength, time-domain coding, spatiotemporal coding.
@@ I Introduction
Since URLLC relies on shorter packet length and smaller transmission interval [6], the length of codewords will be substantially reduced to meet the stringent delay constraint. Finite blocklength coding addresses the practical latency constraints by optimizing the trade-off between blocklength (latency), decoding error probability (reliability), and coding rate [7]. In [8] and [9], the nonasymptotic upper and lower bounds and normal approximations on coding rate as a function of decoding error probability and blocklength at finite blocklength for the binary erasure channel (BEC), the binary symmetric channel (BSC), and the additive white Gaussian noise (AWGN) channel are summarized and analyzed. In a succession of subsequent research efforts, these results have been extended to other types of scalar channels, such as [10, 11, 12, 13]. Obviously, in scalar channels, the amount of data information decreases with decreasing blocklength for given decoding error probability, and it evidently fails to meet the progressively increasing demands of 6G URLLC.
@@
Multiple-input multiple-output (MIMO) technology introduces additional spatial domain degree of freedom (DoF) by deploying multiple antennas at both ends of the transceivers [14]. MIMO is considered to be a key technology for guaranteeing communication performance under low latency requirements in URLLC scenarios [15, 16, 17]. Despite the fact that the normal approximation is not as accurate as the saddlepoint method [18, 19, 20] at low coding rates, its concise form makes it easier to intuitively analyze the relationship between various KPIs and theoretically illustrate the internal logic of MIMO supporting URLLC transmission. There have been a series of research efforts to extend the normal approximation in [8] to MIMO systems. In [21] and [22], the coding rate of a quasi-static MIMO channel was described given blocklength and decoding error probability. [23] presented the expression of the channel dispersion in MIMO coherent fading channel, and [24] presented a high-SNR normal approximation of the maximum coding rate in a non-coherent Rayleigh block-fading MIMO channel. Given that the rate-latency-reliability relationship obtained from the aforementioned research is overly complicated to directly infer the effect of the number of antennas on the performance of a finite blocklength MIMO system, [25] further derived the explicit solutions for the average maximal achievable rate in MIMO systems with respect to spatial DoF, decoding error probability and blocklength based on the conclusion of [21]. The results show that the spatial DoF (coding in spatial-temporal domain) can effectively compensate for the performance loss caused by the reduction of temporal DoF due to shortened codewords. [26] provided the theoretical upper bound for spatiotemporal two-dimensional (2-D) channel coding and suggested that through spatiotemporal coding, transmission reliability and latency can be flexibly balanced in a variety of communication scenarios including URLLC. We can find that the above finite blocklength theoretical studies on MIMO channels all adopt the joint spatiotemporal coding scheme with multiple antennas at the transmitter. However, the time-domain coding scheme with independent coding at each transmit antenna can also be used in practical MIMO coding. A spatiotemporal 2-D polar coding scheme was proposed in [27], and simulation analysis reveals that spatiotemporal coding is capable of fully exploiting the spatial domain resources to achieve high data-rate performance under low latency constraint, while time-domain coding will not benefit from spatial resources in comparison.
@@
To the best of our knowledge, the intuitive performance superiority of spatiotemporal coding over time-domain coding has not yet been thoroughly investigated from a theoretical point of view. In order to comprehensively understand the benefits of spatiotemporal coding for URLLC transmission in MIMO systems, it is essential to derive explicit solutions for the achievable rates of MIMO systems employing spatiotemporal and time-domain coding schemes. The theoretical tools in [8, 9, 21, 22, 23, 24, 25, 26] enable one to characterize the maximal achievable rates of spatiotemporal coding and time-domain coding by analytical means, and it is of great significance to additionally find an explicit expression to identify how spatial DoF improves overall performance. The novel theoretical studies in finite blocklength coding have the potential to provide a new perspective into practical spatiotemporal coding design.
@@
The remainder of the paper is organized as follows. In Section II, we give some preliminaries. The explicit performance bounds of finite blocklength coded MIMO in different coding modes are derived in Section III. Section IV analyzes the normalized maximal achievable rate and decoding error probability. Section V provides numerical results to reveal the advantages of spatiotemporal coding versus time-domain coding in finite blocklength coded MIMO.
@@ II Preliminaries
In this section, we introduce the definitions of information density and binary hypothesis testing under the channel model fixing the output distribution (such as AWGN channel). Based on this, we provide the achievability bound and converse bound to facilitate the subsequent solution of the normal approximations for different coding schemes of the MIMO channel.
@@ III Performance Bounds of MIMO Systems under Different Coding Modes
In this section, we first present the MIMO channel model and analyze the difference between spatiotemporal coding and time-domain coding. Subsequently, we proceed to solve the performance bounds that MIMO channels can achieve in spatiotemporal coding and time-domain coding, respectively.
@@ III-A MIMO Channel Model
We consider a memoryless quasi-static flat Rayleigh fading MIMO channel, wherein the random fading coefficients remain invariant throughout the duration of each codeword. This is a typical assumption for URLLC where the blocklength is usually short enough. The relationship between channel input and channel output can be expressed as
@@
For channel coding in MIMO systems, two coding methods with different utilization of spatial DoF, spatiotemporal coding and time-domain coding, can be considered. The difference between the two methods is illustrated in Fig. 1. In spatiotemporal coding, each bit stream is jointly encoded within both the time and spatial domain, thereby generating a complete codeword $\mathsf{X}$. Then, the corresponding signals of the codeword in different spatial domains are sent out through $L$ transmit antennas. In time domain coding, $L$ bit streams are encoded in the time domain. The codewords ${{\mathbf{x}}_{1}},\ldots,{{\mathbf{x}}_{L}}$ corresponding to each column of the resulting transmitted signal $\mathsf{X}$ are independent of each other. Thereafter, the time-domain codewords are dispatched through $L$ transmit antennas.
@@ III-B Spatiotemporal Coding
$=n\log\det\left({\mathsf{I}_{N}}+\frac{1}{n}{\mathsf{H}^{\text{H}}}\mathsf{Q}_{X}\mathsf{H}\right)-\text{tr}\left[{{\mathbb{W}}^{\text{H}}}\mathbb{W}\right]\log e$
$\quad+\text{tr}\left[{{{\left(\mathsf{X}\mathsf{H}+\mathbb{W}\right)}^{\text{H}}}\left(\mathsf{X}\mathsf{H}+\mathbb{W}\right)}\left({{\mathsf{I}_{N}}+\frac{1}{n}{\mathsf{H}^{\text{H}}}\mathsf{Q}_{X}\mathsf{H}}\right)^{-1}\right]\log e$
$=n\log\det\left({\mathsf{I}_{N}}+\frac{1}{n}{\mathsf{H}^{\text{H}}}\mathsf{Q}_{X}\mathsf{H}\right)+\text{tr}\bigg{[}\left({{\mathsf{I}_{N}}+\frac{1}{n}{\mathsf{H}^{\text{H}}}\mathsf{Q}_{X}\mathsf{H}}\right)^{-1}$
$\quad\left({-\frac{1}{n}{\mathsf{H}^{\text{H}}}\mathsf{Q}_{X}\mathsf{H}{{\mathbb{W}}^{\text{H}}}\mathbb{W}+{{\mathbb{W}}^{\text{H}}}\mathsf{X}\mathsf{H}+{\mathsf{H}^{\text{H}}}{\mathsf{X}^{\text{H}}}\mathbb{W}+\mathsf{H}^{\text{H}}{\mathsf{X}^{\text{H}}}\mathsf{X}\mathsf{H}}\right)\bigg{]}\log{e},$ (17)
@@
Due to the spherical symmetry of both $\mathcal{F}^{\text{ST}}$ (equivalent to ${{P}_{{\mathbb{Y}}|{\mathbb{X}=\mathsf{X},\mathbb{H}=\mathsf{H}}}}$) and capacity achieving output distribution
@@
$\iota\left(\mathsf{X};\mathbb{Y}|\mathsf{H}\right)$
$=n\log\det\left({\mathsf{I}_{N}}+\frac{\rho}{L}{\mathsf{H}^{\text{H}}}\mathsf{H}\right)+\text{tr}\bigg{[}\left({{\mathsf{I}_{N}}+\frac{\rho}{L}{\mathsf{H}^{\text{H}}}\mathsf{H}}\right)^{-1}$
$\qquad\left({-\frac{\rho}{L}{\mathsf{H}^{\text{H}}}\mathsf{H}{{\mathbb{W}}^{\text{H}}}\mathbb{W}+{{\mathbb{W}}^{\text{H}}}\mathsf{X}\mathsf{H}+{\mathsf{H}^{\text{H}}}{\mathsf{X}^{\text{H}}}\mathbb{W}+n\frac{\rho}{L}\mathsf{H}^{\text{H}}\mathsf{H}}\right)\bigg{]}\log{e}$
@@
$\quad+\text{tr}\bigg{[}n{\mathsf{I}_{N}}-\left({\frac{\rho}{L}{\mathsf{H}^{\text{H}}}\mathsf{H}{{\mathbb{W}}^{\text{H}}}\mathbb{W}-{{\mathbb{W}}^{\text{H}}}\mathsf{XH}-{\mathsf{H}^{\text{H}}}{\mathsf{X}^{\text{H}}}\mathbb{W}+n{\mathsf{I}_{N}}}\right)$
$\qquad\qquad\qquad\left({{\mathsf{I}_{N}}+\frac{\rho}{L}{\mathsf{H}^{\text{H}}}\mathsf{H}}\right)^{-1}\bigg{]}\log e$
@@
where ${{W}_{i,j}}\sim\mathcal{CN}\left(0,1\right)$.
Proposition 1
Using the Berry-Esseen CLT, the Neyman-Pearson function for the MIMO channel under spatiotemporal coding can be expressed as
${{\beta}_{\alpha}^{n}}\left({{P}_{{\mathbb{Y}}|{\mathbb{X}=\mathsf{X},\mathbb{H}=\mathsf{H}}}},{{P}_{{\mathbb{Y}}|\mathbb{H}=\mathsf{H}}}\right)$
$\qquad=-n{{C}^{\text{ST}}}\left(\mathsf{H}\right)+\sqrt{n{{V}^{\text{ST}}}\left(\mathsf{H}\right)}\Phi^{-1}\left(1-\alpha\right)+\mathcal{O}\left(\log n\right),$ (19)
where the channel capacity and channel dispersion of the MIMO channel with spatiotemporal codewords in a channel use are expressed as
@@
${{V}^{\text{ST}}}\left(\mathsf{H}\right)$ $=\frac{1}{n}\text{Var}\left[\iota\left(\mathsf{X};\mathbb{Y}|\mathsf{H}\right)\right]$
$=\sum\limits_{i=1}^{m}{\left[1-\frac{1}{{{\left(1+\frac{\rho}{L}{{\lambda}_{i}}\right)}^{2}}}\right]}{{\log}^{2}}e=\sum\limits_{i=1}^{m}{{{V}_{i}}}.$ (21)
We define that